# Half canonical rings of Gorenstein spin curves of genus two

### v1

@@ 2 $\frac{(1-t^{2})(1-t^{12})}{(1-t)(1-t)(1-t^{4})(1-t^{6})}$
$\ H^{0}(K_{C})\operatorname{\otimes}H^{0}\left({\mathcal{L}}\left((m-1)\right)K_{C}\right)\to H^{0}({\mathcal{L}}(mK_{C}))=R_{2m+1}$
@@
is surjective for all $m\geq 3$. It follows that $R$ is generated in degrees $\leq 10$. Since the canonical ring is generated in degrees $\leq 6$ (Proposition 3.3), we also know that there are no new generators in degree $10$square
@@ 5.C.1. Type $A(1)$
A basis of in $R_{12}$ is $x^{12},x^{10}y,\dots,y^{6},x^{6}u,x^{4}yu,x^{2}y^{2}u,y^{3}u$, so comparing the coefficients of the monomials involving $u$ we see that $\lambda=0$, $f_{4}=0$, $xz=u$ and we have the claimed relation $z^{2}-f_{10}=0$. Comparing the Hilbert series (cf. Proposition 5.5), we see that we have found the full ring.
@@ 5.C.2. Type $A(0)$
Comparing the Hilbert series (cf. Proposition 5.5) we see that there are no additional relations. Assume that $\alpha=\beta=\gamma=0$: then $C$ contains the point $q=[0,0,0,0,1]$ and for a general choice of $a_{6},b_{6},c_{6}$ the tangent cone to $C$ at this point has six irreducible components, so $C$ does not have an $A_{n}$ singularity at $q$, contradicting Corollary 3.9. Therefore at least one of $\alpha,\beta,\gamma$ is nonzero and we can use one of the relations above to eliminate $u$. Up to a linear change of the coordinates $z_{1},z_{2}$ the remaining relations can be written either as:
@@ 5.D. Type B: talking ’bout an involution
By Proposition 4.3 there are a partial normalisation $f\colon\widetilde{C}\to C$ and a regular ggs structure (~$(\widetilde{\mathcal{L}},\widetilde{\mu})$ on $\widetilde{C}$ such that ${\mathcal{L}}=f_{*}\widetilde{\mathcal{L}}$.
@@
$(ii)$ Linearisations of ${\mathcal{L}}$, $\widetilde{\mathcal{L}}$ with respect to $\iota$, $\widetilde{\iota}$ can be seen to be in bijection arguing as in the proof of Lemma 2.3, $(ii)$. By relative duality there is a ${\mathcal{O}}_{C}$-module homomorphism $f_{*}\omega_{\widetilde{C}}\to\omega_{C}$ which is invariant with respect to the natural linearisations of these sheaves with respect to $\widetilde{\iota}$ and $\iota$ (cf. Remark 3.7). So it is enough to show that for any given linearisation of $\widetilde{\mathcal{L}}$, the isomorphism ~~$\widetilde{\mathcal{L}}\otimes\widetilde{\mathcal{L}}\to\omega_{\widetilde{C}}$ induces on $\omega_{\widetilde{C}}$ the opposite of the natural linearisation.
@@
By [FFP22, Lemma 4.4] there is a trivializing open cover $\{U_{j}\}$ for $\widetilde{\mathcal{L}}$ on $\widetilde{C}$ such that the $U_{j}$ are $\widetilde{\iota}$-invariant. Let $\sigma_{j}$ be a local generator of ~$\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{U_{j}}$}}$ and let $q\in U_{j}$ be a point: then at least one between $\sigma_{j}+{\widetilde{\iota}}^{*}\sigma_{j}$ and $\sigma_{j}-{\widetilde{\iota}}^{*}\sigma_{j}$ does not vanish at $q$. So, up to refining the cover, we may assume in addition that for every $j$ the sheaf ~$\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{U_{j}}$}}$ is generated by an element $\tau_{j}$ such that ${\widetilde{\iota}}^{*}\tau_{j}=\pm\tau_{j}$. Via the isomorphism ~~$\widetilde{\mathcal{L}}\operatorname{\otimes}\widetilde{\mathcal{L}}\to\omega_{\widetilde{C}}$ we obtain a local trivialization of $\omega_{\widetilde{C}}$ such that for every $j$ there is a local generator $\alpha_{j}$ on $U_{j}$ that verifies ${\widetilde{\iota}}^{*}\alpha_{j}=\alpha_{j}$. In other words, $\omega_{\widetilde{C}}$ is the pull back of a line bundle $\widetilde{\mathcal{M}}$ on $\widetilde{C}/\widetilde{\iota}$ and the linearisation induced by the ggs structure coincides with the linearisation as the pullback of $\widetilde{\mathcal{M}}$. By Proposition 4.5 ${\mathcal{L}}$ is not regular at the intersection point of $C_{1}$ and $C_{2}$, hence $\widetilde{C}=\widetilde{C}_{1}\sqcup\widetilde{C}_{2}$ is a disjoint union of curves of arithmetic genus 0 or 1. The bicanonical involution $\iota$ induces an involution $\widetilde{\iota}$ of $\widetilde{C}$ that preserves the two components and such that $Q_{i}:=\widetilde{C}_{i}/\widetilde{\iota}$ is smooth rational for $i=1,2$. So the quotient map $\varphi_{i}\colon\widetilde{C}_{i}\to Q_{i}$ is flat and ${\varphi_{i}}_{*}{\mathcal{O}}_{\widetilde{C}_{i}}={\mathcal{O}}_{Q_{i}}\oplus{\mathcal{O}}_{Q_{i}}(-a)$, where $a=p_{a}(C_{i})+1$. Relative duality gives ${\varphi_{i}}_{*}\omega_{\widetilde{C}_{i}}=\omega_{Q_{i}}\oplus\omega_{Q_{i}}(a)$. Since $\deg\omega_{\widetilde{C}_{i}}=2\deg{\mathcal{M}}{\raisebox{-0.1pt}{$|_{Q_{i}}$}}$, we conclude that ${\mathcal{M}}{\raisebox{-0.1pt}{$|_{Q_{i}}$}}=\omega_{Q_{i}}(a)$, namely the linearisation induced on $\omega_{C_{i}}$ by the isomorphism ~~$\widetilde{\mathcal{L}}\operatorname{\otimes}\widetilde{\mathcal{L}}\to\omega_{\widetilde{C}}$ is the opposite of the natural one. $\square$
@@
Since ${\mathcal{L}}$ is not locally free at the intersection point $p$ of $C_{1}$ and $C_{2}$ (cf. Proposition 4.5), by Proposition 4.3 there is a partial normalisation $f\colon\widetilde{C}_{1}\sqcup\widetilde{C}_{2}\to C$ such that ${\mathcal{L}}=f_{*}\widetilde{\mathcal{L}}$, where ~=~$\widetilde{\mathcal{L}}=\widetilde{\mathcal{L}}_{1}\oplus\widetilde{\mathcal{L}}_{2}$ and ~$\widetilde{\mathcal{L}}_{i}$ is a line bundle supported on $\widetilde{C}_{i}$. Recall that $\widetilde{C}_{i}$ is Gorenstein, $p_{a}(\widetilde{C}_{i})=0$ or 1 and ~$\widetilde{\mathcal{L}}_{i}^{\otimes 2}=\omega_{\widetilde{C}_{i}}$. We set ${\mathcal{L}}_{i}:=f_{*}\widetilde{\mathcal{L}}_{i}$ and $a_{i}:=h^{0}({\mathcal{L}}_{i})=h^{0}(\widetilde{\mathcal{L}}_{i})$. We have $a_{1}+a_{2}=h^{0}({\mathcal{L}})\leq 2$ by Proposition 4.7, so if we assume $a_{1}\leq a_{2}$ the possibilities for the pair $(a_{1},a_{2})$ are $(0,0)$, $(0,1)$ or $(1,1)$.
@@
${\varphi_{i}}_{*}{{\mathcal{L}}_{i}(np_{i})}=\begin{cases}({\varphi_{i}}_{*}{\mathcal{L}}_{i})\otimes{\mathcal{O}}_{Q_{i}}(\tfrac{n}{2})&\text{ $n$ even,}\\ ({\varphi_{i}}_{*}{\mathcal{L}}_{i}(p_{i}))\otimes{\mathcal{O}}_{Q_{i}}(\tfrac{n-1}{2})&\text{ $n$ odd.}\end{cases}$
@@
The line bundle ~$\widetilde{\mathcal{L}}_{i}(p_{i})$ has degree $p_{a}(\widetilde{C}_{i})$. If $p_{a}(\widetilde{C}_{i})=0$, then ~$\widetilde{\mathcal{L}}_{i}(p_{i})={\mathcal{O}}_{\widetilde{C}_{i}}$ and ${\varphi_{i}}_{*}\widetilde{\mathcal{L}}_{i}(p_{i})={\mathcal{O}}_{Q_{i}}\oplus{\mathcal{O}}_{Q_{i}}(-1)$. If $p_{a}(\widetilde{C}_{i})=1$, then $h^{0}(\widetilde{\mathcal{L}}_{i}(p_{i}))=1$ and $h^{1}(\widetilde{\mathcal{L}}_{i}(p_{i}))=0$, so that ${\varphi_{i}}_{*}\widetilde{\mathcal{L}}_{i}(p_{i})={\mathcal{O}}_{Q_{i}}\oplus{\mathcal{O}}_{Q_{i}}(-1)$ also in this case. So for $i=1,2$ we choose the linearization of ~$\widetilde{\mathcal{L}}_{i}(p_{i})$ in such a way that ${\mathcal{O}}_{Q_{i}}$ is the invariant summand; this is the same as choosing a linearisation of ${\mathcal{L}}\otimes\omega_{C}$ as claimed in $(i)$. Note that choosing a linearisation of ${\mathcal{L}}\otimes\omega_{C}$ is the same as choosing a linearisation of ${\mathcal{L}}$, because $\omega_{C}$ is invertible and has a chosen linearisation (cf. Remark 3.7 and Lemma 5.9). Note also that if $H^{0}({\mathcal{L}})$ contains a nonzero section $x$, then $x^{3}\in H^{0}({\mathcal{L}}\otimes\omega_{C})$ is invariant, hence so is $x$ and claim (i) is proven. Statement $(ii)$ for odd $n>0$ follows from the above remarks.
@@ 5.D.1. Type $B(1)$
2×2 minors of ($2\times 2\text{ minors of }\begin{pmatrix}0&y&w&z\\ x&w&v&u\end{pmatrix}\ \text{ and }\begin{array}[]{rcl}z^{2}&=&yg_{8}(y,v)\\ zu&=&wg_{8}(y,v)\\ u^{2}&=&vg_{8}(y,v)+x^{4}h_{8}(x,v)\end{array}$
@@
$R(C,K_{C})\cong{\mathbb{C}}[x^{2},y,v,u]/(x^{2}y,u^{2}-f_{12}(x^{2},y,v))\supset R(C,K_{C})^{+}={\mathbb{C}}[x^{2},y,v]/(x^{2}y).$
@@
$R_{5}^{+}\cong H^{0}({{\mathcal{O}}_{Q_{1}}})\oplus H^{0}({{\mathcal{O}}_{Q_{2}}}(1))\text{ and }R_{5}^{-}\cong H^{0}({\mathcal{O}}_{Q_{1}})\oplus H^{0}({\mathcal{O}}_{Q_{2}}(-1)).$
@@
We have analysed the situation up to degree $5$ already, so we start in degree $6$. By Proposition 3.3, $R_{6}$ is generated by $x^{2},y,v,u$ and $u$ does not vanish at the generic points of $C_{1}$ and $C_{2}$. Clearly, $w^{2}$ is an invariant section which is supported on $C_{1}$, so we get a relation
@@
thus there are no new relations in degree $7$. Indeed, consider the first summand of $R_{7}^{+}$ and assume that there is a linear relation $\alpha wy^{2}+\beta wv=0$. At the generic point of $C_{1}$ we can divide by $w$ to get $\alpha y^{2}+\beta v=0$ on $C_{1}$, namely $\alpha y^{2}+\beta v=\gamma x^{4}$ for some scalar $\gamma$. Since $y^{2},x^{4},v$ are a basis of $R_{4}$ we get $\alpha=\beta=\gamma=0$ and $wy^{2},wv$ are independent. We can argue in a similar way for the second summand of $R_{7}$, dividing by $x^{3}$. The two summands of $R_{7}^{-}$ are easily treated by noting that $xu\neq 0$ and $yz\neq 0$.
@@
We proceed just as we did in degree $7$. At the generic point of $C_{1}$, we can divide by $w$ to show that $wyv$ and $wy^{3}$ are linearly independent. Hence the first summand of $R_{9}^{+}$ is as claimed. For the second summand, we can divide by $x$ at the generic point of $C_{2}$.

### v2

@@ 2 $\frac{(1-t^{2})(1-t^{12})}{(1-t)(1-t)(1-t^{4})(1-t^{6})}$
$H^{0}(K_{C})\operatorname{\otimes}H^{0}\left({\mathcal{L}}((m-1)K_{C})\right)\to H^{0}({\mathcal{L}}(mK_{C}))=R_{2m+1}$
@@
is surjective for all $m\geq 3$. It follows that $R$ is generated in degrees $\leq 10$. Since the canonical ring is generated in degrees $\leq 6$ (Proposition 3.3), we also know that there are no new generators in degree $10$. $\square$
@@ 5.C.1. Type $A(1)$
A basis of $R_{12}$ is $x^{12},x^{10}y,\dots,y^{6},x^{6}u,x^{4}yu,x^{2}y^{2}u,y^{3}u$, so comparing the coefficients of the monomials involving $u$ we see that $\lambda=0$, $f_{4}=0$, $xz=u$ and we have the claimed relation $z^{2}-f_{10}=0$. Comparing the Hilbert series (cf. Proposition 5.5), we see that we have found the full ring.
@@ 5.C.2. Type $A(0)$
Comparing the Hilbert series (cf. Proposition 5.5) we see that there are no additional relations and we can embedd $C$ into ${\mathbb{P}}(2,2,3,3,6)$. Assume that $\alpha=\beta=\gamma=0$: then $C$ contains the point $q=[0,0,0,0,1]$ and for a general choice of $a_{6},b_{6},c_{6}$ the tangent cone to $C$ at this point has six irreducible components since the 3 equations give $6$ points in $\mathbb{P}(2,2,3,3)$ and the curve is the cone over these 6 points with vertex $q$, so $C$ does not have an $A_{n}$ singularity at $q$, contradicting Corollary 3.9. Therefore at least one of $\alpha,\beta,\gamma$ is nonzero and we can use one of the relations above to eliminate $u$. Up to a linear change of the coordinates $z_{1},z_{2}$ the remaining relations can be written either as:
@@ 5.D. Type B: talking ’bout an involution
[…]
@@
$(ii)$ Linearisations of $\mathcal{L}$}, $\widetilde{\mathcal{L}}$ with respect to $\iota$, $\widetilde{\iota}$ can be seen to be in bijection arguing as in the proof of Lemma 2.4, $(ii)$. By relative duality there is a ${\mathcal{O}}_{C}$-module homomorphism $f_{*}\omega_{\widetilde{C}}\to\omega_{C}$ which is invariant with respect to the natural linearisations of these sheaves with respect to $\widetilde{\iota}$ and $\iota$ (cf. Remark 3.7). So it is enough to show that for any given linearisation of $\widetilde{\mathcal{L}}$, the isomorphism $\widetilde{\mathcal{L}}\otimes\widetilde{\mathcal{L}}\to\omega_{\widetilde{C}}$ induces on $\omega_{\widetilde{C}}$ the opposite of the natural linearisation.
@@
By [FFP22, Lemma 4.4] there is a trivializing open cover $\{U_{j}\}$ for $\widetilde{\mathcal{L}}$ on $\widetilde{C}$ such that the $U_{j}$ are $\widetilde{\iota}-$invariant. Let $\sigma_{j}$ be a local generator of $\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{U_{j}}$}}$ and let $q\in U_{j}$ be a point: then at least one between $\sigma_{j}+{\widetilde{\iota}}^{*}\sigma_{j}$ and $\sigma_{j}-{\widetilde{\iota}}^{*}\sigma_{j}$ does not vanish at $q$. So, up to refining the cover, we may assume in addition that for every $j$ the sheaf $\widetilde{\mathcal{L}}{\raisebox{-0.1pt}{$|_{U_{j}}$}}$ is generated by an element $\tau_{j}$ such that ${\widetilde{\iota}}^{*}\tau_{j}=\pm\tau_{j}$. Via the isomorphism $\widetilde{\mathcal{L}}\operatorname{\otimes}\widetilde{\mathcal{L}}\to\omega_{\widetilde{C}}$ we obtain a local trivialization of $\omega_{\widetilde{C}}$ such that for every $j$ there is a local generator $\alpha_{j}$ on $U_{j}$ that verifies ${\widetilde{\iota}}^{*}\alpha_{j}=\alpha_{j}$. In other words, $\omega_{\widetilde{C}}$ is the pull back of a line bundle $\widetilde{\mathcal{M}}$ on $\widetilde{C}/\widetilde{\iota}$ and the linearisation induced by the ggs structure coincides with the linearisation as the pullback of $\widetilde{\mathcal{M}}$. By Proposition 4.4 $\mathcal{L}$} is not regular at the intersection point of $C_{1}$ and $C_{2}$, hence $\widetilde{C}=\widetilde{C}_{1}\sqcup\widetilde{C}_{2}$ is a disjoint union of curves of arithmetic genus 0 or 1. The bicanonical involution $\iota$ induces an involution $\widetilde{\iota}$ of $\widetilde{C}$ that preserves the two components and such that $Q_{i}:=\widetilde{C}_{i}/\widetilde{\iota}$ is smooth rational for $i=1,2$. So the quotient map $\varphi_{i}\colon\widetilde{C}_{i}\to Q_{i}$ is flat and ${\varphi_{i}}_{*}{\mathcal{O}}_{\widetilde{C}_{i}}={\mathcal{O}}_{Q_{i}}\oplus{\mathcal{O}}_{Q_{i}}(-a)$, where $a=p_{a}(C_{i})+1$. Relative duality gives ${\varphi_{i}}_{*}\omega_{\widetilde{C}_{i}}=\omega_{Q_{i}}\oplus\omega_{Q_{i}}(a)$. Since $\deg\omega_{\widetilde{C}_{i}}=2\deg{\mathcal{M}}{\raisebox{-0.1pt}{$|_{Q_{i}}$}}$, we conclude that ${\mathcal{M}}{\raisebox{-0.1pt}{$|_{Q_{i}}$}}=\omega_{Q_{i}}(a)$, namely the linearisation induced on $\omega_{C_{i}}$ by the isomorphism $\widetilde{\mathcal{L}}\operatorname{\otimes}\widetilde{\mathcal{L}}\to\omega_{\widetilde{C}}$ is the opposite of the natural one. $\square$
@@
Since $\mathcal{L}$} is not locally free at the intersection point $p$ of $C_{1}$ and $C_{2}$ (cf. Proposition 4.4), by Proposition 4.2 there is a partial normalisation $f\colon\widetilde{C}_{1}\sqcup\widetilde{C}_{2}\to C$ such that ${\mathcal{L}}=f_{*}\widetilde{\mathcal{L}}$, where $\widetilde{\mathcal{L}}=\widetilde{\mathcal{L}}_{1}\oplus\widetilde{\mathcal{L}}_{2}$ and $\widetilde{\mathcal{L}}_{i}$ is a line bundle supported on $\widetilde{C}_{i}$. Recall that $\widetilde{C}_{i}$ is Gorenstein, $p_{a}(\widetilde{C}_{i})=0$ or 1 and $\widetilde{\mathcal{L}}_{i}^{\otimes 2}=\omega_{\widetilde{C}_{i}}$. We set ${\mathcal{L}}_{i}:=f_{*}\widetilde{\mathcal{L}}_{i}$ and $a_{i}:=h^{0}({\mathcal{L}}_{i})=h^{0}(\widetilde{\mathcal{L}}_{i})$. We have $a_{1}+a_{2}=h^{0}({\mathcal{L}})\leq 2$ by Proposition 4.6, so if we assume $a_{1}\leq a_{2}$ the possibilities for the pair $(a_{1},a_{2})$ are $(0,0)$, $(0,1)$ or $(1,1)$.
@@
${\varphi_{i}}_{*}{\widetilde{\mathcal{L}}_{i}(np_{i})}=\begin{cases}({\varphi_{i}}_{*}\widetilde{\mathcal{L}}_{i})\otimes{\mathcal{O}}_{Q_{i}}(\tfrac{n}{2})&\text{ $n$ even,}\\ ({\varphi_{i}}_{*}\widetilde{\mathcal{L}}_{i}(p_{i}))\otimes{\mathcal{O}}_{Q_{i}}(\tfrac{n-1}{2})&\text{ $n$ odd.}\end{cases}$
@@
The line bundle $\widetilde{\mathcal{L}}_{i}(p_{i})$ has degree $p_{a}(\widetilde{C}_{i})$. If $p_{a}(\widetilde{C}_{i})=0$, then $\widetilde{\mathcal{L}}_{i}(p_{i})={\mathcal{O}}_{\widetilde{C}_{i}}$ and ${\varphi_{i}}_{*}\widetilde{\mathcal{L}}_{i}(p_{i})={\mathcal{O}}_{Q_{i}}\oplus{\mathcal{O}}_{Q_{i}}(-1)$. If $p_{a}(\widetilde{C}_{i})=1$, then $h^{0}(\widetilde{\mathcal{L}}_{i}(p_{i}))=1$ and $h^{1}(\widetilde{\mathcal{L}}_{i}(p_{i}))=0$, so that ${\varphi_{i}}_{*}\widetilde{\mathcal{L}}_{i}(p_{i})={\mathcal{O}}_{Q_{i}}\oplus{\mathcal{O}}_{Q_{i}}(-1)$ also in this case. So for $i=1,2$ we choose the linearization of $\widetilde{\mathcal{L}}_{i}(p_{i})$ in such a way that ${\mathcal{O}}_{Q_{i}}$ is the invariant summand; this is the same as choosing a linearisation of ${\mathcal{L}}\otimes\omega_{C}$ as claimed in $(i)$. Note that choosing a linearisation of ${\mathcal{L}}\otimes\omega_{C}$ is the same as choosing a linearisation of $\mathcal{L}$}, because $\omega_{C}$ is invertible and has a chosen linearisation (cf. Remark 3.7 and Lemma 5.9). Note also that if $H^{0}({\mathcal{L}})$ contains a nonzero section $x$, then $x^{3}\in H^{0}({\mathcal{L}}\otimes\omega_{C})$ is invariant, hence so is $x$ and claim (i) is proven. Statement $(ii)$ for odd $n>0$ follows from the above remarks.
@@ 5.D.1. Type $B(1)$
$2\times 2\text{ minors of }\begin{pmatrix}0&y&w&z\\ x&w&v&u\end{pmatrix}\ \text{ and }\begin{array}[]{rcl}z^{2}&=&yg_{8}(y,v)\\ zu&=&wg_{8}(y,v)\\ u^{2}&=&vg_{8}(y,v)+x^{4}h_{8}(x,v)\end{array}$
@@
(5.11) $R(C,K_{C})\cong{\mathbb{C}}[x^{2},y,v,u]/(x^{2}y,u^{2}-f_{12}(x^{2},y,v))\supset R(C,K_{C})^{+}={\mathbb{C}}[x^{2},y,v]/(x^{2}y).$
@@
[…]
@@
We have analysed the situation up to degree $5$ already, so we start in degree $6$. By (5.11), we have
$R_{6}^{+}=\langle x^{2},y,v\rangle\text{ and }R_{6}^{-}=\langle u\rangle$
and $u$ is an anti-invariant section that does not vanish at the generic points of $C_{1}$ and $C_{2}$. Clearly $w^{2}$ is an invariant section which is supported on $C_{1}$, so we get a relation
@@
thus there are no new relations in degree $7$. Indeed, consider the first summand of $R_{7}^{+}$ and assume that there is a linear relation $\alpha wy^{2}+\beta wv=0$. At the generic point of $C_{1}$ we can divide by $w$ to get $\alpha y^{2}+\beta v=0$ on $C_{1}$, namely $\alpha y^{2}+\beta v=\gamma x^{4}$ for some scalar $\gamma$. Since $y^{2},x^{4},v$ are a basis of $R_{4}$ we get $\alpha=\beta=\gamma=0$ and $wy^{2},wv$ are independent. We can argue in a similar way for the second summand of $R_{7}$, dividing by $x^{3}$. The two summands of $R_{7}^{-}$ are easily treated by noting that $xu\neq 0$ and $yz\neq 0$. In this way we have proved the independence of the generators and we can conclude by a dimension count.
@@
We proceed just as we did in degree $7$. At the generic point of $C_{1}$, we can divide by $w$ to show that wyvwyv and $wy^{3}$ are linearly independent. Hence the first summand of $R_{9}^{+}$ is as claimed. For the second summand, we can divide by $x$ at the generic point of $C_{2}$.